# Near-real-time flood inundation monitoring by Bayesian analysis for change point problems for Sentinel-1 time series

### v2

@@ 6.1 Methodological Advances in SAR-Based Flood Detection
Our study introduces Bayesian analysis for change point problems (BCP) to near real-time flood monitoring applied to Sentinel-1 GRD time series data. Unlike supervised methods that require extensive training data and face challenges with geographical transferability [Wu2023-kc, Ghosh2024-ub], our approach leverages the statistical properties of SAR time series to detect flood-induced changes without any training requirements. This fundamental training-free characteristic is a significant advantage, as it completely eliminates the overfitting issues inherent in deep learning and other machine learning approaches. Machine learning methods often perform well in areas represented in training datasets but fail when applied to novel geographical contexts or unusual flood patterns. By contrast, our BCP method requires no supervised learning process, making it immediately applicable to any region with sufficient SAR time series data, regardless of whether similar flood events have been previously observed and labeled.
Our method advances the field by reframing flood detection as a temporal changepoint identification problem rather than a binary classification task. Conventional classification approaches typically rely on absolute backscatter values or simple change metrics between two images, making them vulnerable to variations in instrument calibration, incidence angle effects, and seasonal changes in land cover [Schlaffer2015-uv]. In contrast, our Bayesian framework models the entire temporal trajectory of backscatter values, enabling more robust discrimination between normal fluctuations and flood-induced changes.
Parameter sensitivity analysis showed that a window size of 9 pixels provided a good balance between noise reduction and preservation of flood boundary details, though our method demonstrated relatively stable performance across a range of parameter settings.
@@ 6.2 Performance in Relation to State-of-the-Art Approaches
Our performance analysis reveals distinct patterns across different environmental contexts and detection methods. In open areas, our BCP method demonstrated robust performance with F1 scores ranging from 0.74 to 0.84, outperforming the OTSU thresholding method (F1: 0.031-0.132). The Siamese CNN showed variable performance in open areas, achieving its best results in Weihui (F1: 0.626) but performing less effectively in Jubba (F1: 0.366) and NovaKakhovka (F1: 0.173). This variability suggests that while deep learning approaches can achieve moderate success in open areas, they may be sensitive to regional differences in landscape characteristics and flood patterns.
The stark performance disparity between open and urban areas can be attributed to fundamental differences in SAR backscatter characteristics. Open areas offer more uniform backscatter patterns in SAR imagery, enabling more reliable flood detection across all methods due to clearer contrast between flooded and non-flooded surfaces. In contrast, urban environments present complex scattering mechanisms due to built infrastructure, leading to increased signal ambiguity. The presence of various materials and structures in urban areas creates multiple reflection pathways, making flood detection particularly challenging. These physical constraints fundamentally limit the effectiveness of current detection methods, regardless of their computational sophistication.
@@ 6.3 Integration with Operational Flood Monitoring Systems
The practical application of our method extends beyond academic research to operational flood monitoring systems. Our approach could complement existing near real-time flood mapping services such as the Copernicus Emergency Management Service (EMS) and NASA’s MODIS Near Real-Time Global Flood Mapping system. These systems currently rely primarily on optical imagery or simple thresholding of SAR data, facing challenges with cloud cover and discrimination between permanent and temporary water features.
@@
The method’s ability to process raw Sentinel-1 GRD data directly also streamlines the operational workflow, reducing the need for extensive preprocessing chains that can introduce delays in time-sensitive emergency response scenarios. Combined with Sentinel-1’s systematic acquisition plan, this could enable near-real-time flood alerts within hours of satellite overpass, a critical improvement over current systems that often require days to deliver actionable information.
@@ 6.4 Limitations and Future Research Directions
Despite the promising results, several limitations of our approach warrant consideration. The method’s performance in open areas, particularly in NovaKakhovka, indicates challenges in distinguishing flood inundation from other changes in agricultural regions. This limitation appears related to the complex temporal patterns of vegetation growth and harvest activities that can create backscatter changes similar to flooding. Future research could address this by incorporating ancillary information on land cover and seasonal agricultural calendars to better contextualize detected changes.
From a methodological perspective, the current BCP implementation assumes normally distributed backscatter values, which may not fully capture the complex statistical properties of SAR data. Alternative Bayesian frameworks that incorporate more sophisticated distributions, such as the gamma distribution commonly used for SAR intensity modeling, could potentially improve detection accuracy. Similarly, hierarchical Bayesian models that explicitly account for spatial context could enhance the method’s robustness to speckle noise without requiring post-processing spatial filtering.
The temporal resolution of Sentinel-1 (6-12 days) represents another limitation for monitoring rapidly evolving flood events [Munasinghe2023-oc]. Future research could explore the integration of our approach with other SAR platforms (e.g., RADARSAT Constellation Mission, ALOS-2) to achieve higher temporal frequency. Alternatively, data fusion techniques that combine SAR with optical or passive microwave observations could provide more continuous monitoring capabilities, leveraging the complementary strengths of different sensor types.
@@ 7 Conclusions
We present a novel approach to near real-time flood monitoring using Bayesian Changepoint Detection (BCP) with Sentinel-1 SAR time series data. Our key contributions include: (1) a training-free framework for automated flood detection, (2) innovative application of Bayesian analysis to SAR time series, and (3) empirical validation across diverse contexts.
As climate change increases flood risks globally, our method contributes to disaster resilience and adaptation efforts through its accessibility, transferability, and operational efficiency.
Our approach can be incorporated into semi-real-time flood monitoring systems due to its automatic processing capabilities. However, a key constraint is the infrequent coverage of Sentinel-1, which creates a bottleneck during urgent disaster response. Further research is needed to demonstrate our approach with higher-frequency SAR time series data.
8 Acknowledgements
This work was funded by NEDO (23200859-0).

### v3

@@ 7.1 Methodological Advances in SAR-Based Flood Detection
Our study introduces Bayesian analysis for change point problems (BCP) to flood monitoring applied to Sentinel-1 GRD time series data. Unlike supervised methods that require extensive labeled training data and face challenges with geographical transferability [16, 17], our approach leverages the statistical properties of SAR time series to detect flood-induced changes without any training requirements. This fundamental training-free characteristic is a substantial advantage, as it completely eliminates the overfitting issues inherent in deep learning and other machine learning approaches. Machine learning methods often perform well in areas represented in training datasets but fail when applied to novel geographical contexts or unusual flood patterns [30, 19]. By contrast, our BCP method requires no supervised learning process, making it immediately applicable to any region with sufficient SAR time series data, regardless of whether similar flood events have been previously observed and labeled.
Our method advances the field by reframing flood detection as a temporal changepoint identification problem rather than a binary classification task. Conventional classification approaches typically rely on absolute backscatter values or simple change metrics between two images, making them vulnerable to variations in instrument calibration, incidence angle effects, and seasonal changes in land cover [14]. In contrast, our Bayesian framework models the entire temporal trajectory of backscatter values, enabling more robust discrimination between normal fluctuations and flood-induced changes.
Parameter sensitivity analysis showed that a window size of 9 pixels provided a good balance between noise reduction and preservation of flood boundary details, though our method demonstrated relatively stable performance across a range of parameter settings. Unlike most previous studies that applied fixed parameter settings for speckle suppression in SAR imagery, we explored various parameter combinations to maximize the effectiveness of BCP outputs. Our analysis revealed that $w=9$ and $t=0.2$ yielded the maximum F1 score for our studies (Figure 5). However, it should be noted that these optimal parameters were determined using the test dataset with a spatial resolution of 20-m provided by UrbanSARFloods. Optimal parameter settings should be determined for the original Sentinel-1 GRD data with a spatial resolution of 10-m for real-world applications.
@@ 7.2 Performance in Relation to State-of-the-Art Approaches
Our performance analysis reveals distinct patterns across different environmental contexts and detection methods. In open areas (Table 4), our BCP method demonstrated robust performance with F1 scores ranging from 0.49 to 0.74, outperforming Otsu’s thresholding approach (F1: 0.02-0.14). The Siamese CNN showed variable performance in open areas, achieving its best results in Weihui (F1: 0.63) but performing less effectively in Jubba (F1: 0.36) and Nova Kakhovka (F1: 0.17). These results suggest that while deep learning approaches can achieve moderate success, the model may suffer from insufficient training to identify flood patterns in unknown landscape characteristics.
The stark performance disparity between open and urban areas can be attributed to fundamental differences in SAR backscatter characteristics [19]. Open areas offer more uniform backscatter patterns in SAR imagery, enabling more reliable flood detection across all methods due to clearer contrast between flooded and non-flooded surfaces. In contrast, urban environments present complex scattering mechanisms due to built infrastructure, leading to increased signal ambiguity [30, 19]. The presence of various materials and structures in urban areas creates multiple reflection pathways, making flood detection particularly challenging. These physical constraints fundamentally limit the effectiveness of current detection methods, regardless of their computational sophistication.
@@ 7.3 Integration with Operational Flood Monitoring Systems
The practical application of our method extends beyond academic research to operational flood monitoring systems. Our approach could complement existing near-real-time flood mapping services such as the Copernicus Emergency Management Service (EMS) [46] and NASA’s MODerate Resolution Imaging Spectroradiometer (MODIS)/Visible Infrared Imaging Radiometer Suite (VIIRS) near-real-time global flood products [47, 48, 49]. These systems currently rely primarily on optical imagery or simple thresholding of SAR data, facing challenges with cloud cover and discrimination between permanent and temporary water features.
@@
The method’s ability to process original Sentinel-1 GRD data directly also streamlines the operational workflow, reducing the need for extensive preprocessing chains that can introduce delays in time-sensitive emergency response scenarios. Combined with Sentinel-1’s systematic acquisition plan, this could enable near-real-time flood alerts within hours of satellite overpass, a critical improvement over current systems that often require days to deliver actionable information.
@@ 7.4 Limitations and Future Research Directions
Despite the promising results, several limitations of our approach warrant consideration. The method’s performance in open areas, particularly in Nova Kakhovka, indicates challenges in distinguishing flood inundation from other changes in agricultural regions. This limitation appears related to the complex temporal patterns of vegetation growth and harvest activities that can create backscatter changes similar to flooding [30]. Additionally, our approach often misclassify non-flood changes in SAR backscatter as floods due to complex scattering mechanisms that impair water detection, particularly in urban environments. [30, 19]. Future research could address these limitations by developing more sophisticated time series analysis techniques that can better discriminate between different types of change patterns within the SAR signal itself, enhancing the algorithm’s ability to identify unique flood signatures without relying on external datasets.
From a methodological perspective, the current BCP implementation assumes normally distributed observation values [23], which may not fully capture the complex statistical properties of SAR data. Alternative Bayesian frameworks that incorporate more sophisticated distributions, such as the gamma distribution commonly used for SAR intensity modeling, could potentially improve detection accuracy. Similarly, hierarchical Bayesian models that explicitly account for spatial context could enhance the method’s robustness to speckle without requiring post-processing spatial filtering. These approaches are specifically designed to address the challenge of flood detection in urban areas.
The temporal resolution of Sentinel-1 (6-12 days) represents another limitation for monitoring rapidly evolving flood events [50]. Future research could explore the integration of our approach with constellation SAR observations to achieve higher temporal frequency. Alternatively, data fusion techniques that combine SAR with optical or passive microwave observations could provide more continuous monitoring capabilities, leveraging the complementary strengths of different sensor types.
@@ 8 Conclusions
We present a novel training-free approach for automated flood detection using Bayesian analysis for change point problems (BCP) applied to Sentinel-1 GRD time series data. Our method analyzes one-year baseline periods of SAR backscatter intensity to statistically model normal temporal behavior, then computes posterior probabilities of change points at flood observation dates. Using combined VV and VH polarizations with optimized post-processing parameters (moving-window kernel size = 9, threshold = 0.2), the approach was validated across three diverse geographical contexts.
The BCP method substantially outperformed baseline approaches, particularly when compared to the Siamese CNN deep learning technique, achieving F1 scores 140–214% higher. A key benefit of our approach is that it requires no training data, unlike deep learning methods. Across all test sites, BCP consistently outperformed baseline methods, with F1 scores of 0.75 (Weihui), 0.68 (Jubba), and 0.49 (Nova Kakhovka).
The training-free nature of our approach eliminates overfitting issues and enables immediate deployment to any region with sufficient SAR coverage, regardless of whether similar flood events have been previously observed. This accessibility is particularly valuable for disaster response in data-scarce regions where technical resources are limited. Our method’s ability to process original Sentinel-1 GRD data directly streamlines operational workflows, potentially enabling near-real-time flood alerts within hours of satellite overpass. However, the 6-12 day temporal resolution of Sentinel-1 remains a constraint for monitoring rapidly evolving flood events. Future research should explore integration with higher-frequency SAR constellations and investigate more sophisticated Bayesian frameworks that better capture the complex statistical properties of SAR data, particularly for improving urban flood detection where current performance remains challenging.
Acknowledgments
We appreciate insightful comments and feedback from anonymous reviewers. This work was funded by the Disaster Prevention Research Institute’s Implementation Science Research for Regional Communities (Specific) project at Kyoto University (Project No. 2024RS-01) and NEDO (23200859-0).